# supplementary

### v1

@@ Abstract
Understanding how signals propagate through neural circuits is central to deciphering brain computation. While functional connectivity methods capture statistical associations, they do not identify directionality or mechanisms of influence. We introduce CITS (Causal Inference in Time Series), a non-parametric algorithm for inferring statistical causal neural circuitry from high-resolution neural time series data. CITS models neural dynamics using a time-invariant structural causal model framework under arbitrary Markov order assumptions and tests for time-lagged conditional independence using either Gaussian or distribution-free statistics. Unlike classical Granger Causality, which often relies on linear autoregressive models and Gaussian noise assumptions, or the Peter-Clark algorithm, which assumes i.i.d. data and lacks temporal structure, CITS is designed for temporally dependent, potentially non-Gaussian data with flexible statistical testing. We prove consistency under mild mixing conditions and validate the method on simulated linear, nonlinear, and continuous-time recurrent neural network data, where CITS outperforms state-of-the-art baselines. We then used CITS to reveal putative neural circuits in mouse brain during visual tasks from a large-scale Neuropixels recordings. CITS uncovers interpretable, stimulus-specific putative neural circuitry linking cortical, thalamic, and hippocampal regions that is consistent with existing experimental literature. Additionally, CITS offers new biological insights about how neurons with similar orientation selectivity indices might be more likely to form a statistically causal connected network. These findings demonstrate the ability of CITS in generating biological hypothesis for future experimental studies. Overall, this study provides a theoretically grounded and empirically validated approach to reveal statistically causal pathways in large-scale neural time series.
@@ 1 Introduction
We evaluate CITS on a suite of synthetic time series benchmarks, including linear Gaussian, nonlinear non-Gaussian, and continuous-time recurrent neural networks. In each setting, CITS consistently recovers the correct causal structure more accurately than classical methods such as Granger Causality, the PC algorithm, and its time-aware extension (TPC), particularly in the presence of nonlinear dependencies and indirect causal effects. We further demonstrate the method’s practical utility in neuroscience by applying it to high resolution electrophysiology recordings from mouse visual brain. The resulting causal graphs reveal biologically interpretable, stimulus-specific interactions among cortical, thalamic, and hippocampal regions, highlighting the potential of CITS to uncover directed neural interactions from large-scale electrophysiological time series. Causally connected neuron pairs exhibit significantly higher tuning similarity, supporting the functional relevance of the inferred pathways.
2 CITS: A Statistically Causal Inference Framework for High-Resolution Neural Time Series
@@
[…]
@@ 2.1 Markovian Structural Causal Model for Time Series
[…]
@@
[…]
@@ 2.3.1 Oracle Version: With Conditional Independence Access
[…]
@@
6    if $X_{u,s}\perp\!\!\!\!\perp X_{v,2\tau+1}\leavevmode\nobreak\ |\leavevmode\nobreak\ \bm{X}_{S}$ then
@@ 2.4 Assigning Edge Weights in the Inferred Graphs
$w_{u,s}^{v,t}=\left\{\begin{array}[]{lr}0,&\text{ if }X_{u,s}\not\in pa_{\hat{G}}(X_{v,t}),\\ &\\ \text{coefficient of }X_{u,s}\text{ in }&\\ X_{v,t}\sim pa_{\hat{G}}(X_{v,t})&\text{ if }X_{u,s}\in pa_{\hat{G}}(X_{v,t})\end{array}\right.$ (4)
@@ Discussion.
This corollary follows from the temporal structure of the model. When all causal effects among observables are strictly time-lagged (i.e., from $X_{u,t-1}$ to $X_{v,t}$), with no concurrent (within-time-slice) effects, and the system is one-step Markovian, latent variables can induce at most contemporaneous dependencies within a single time step. Since CITS only tests for lagged conditional dependencies and conditions on past variables, it avoids inferring spurious lagged edges due to latent confounding under these assumptions. As long as the marginal distribution over observed variables is faithful to the induced causal graph, the consistency guarantees of Theorem 2.2 continue to apply.
The assumption that the marginal distribution over observed variables is faithful to the marginal causal graph is standard in causal discovery. While faithfulness is not guaranteed under marginalization, since conditional independencies can arise through parameter cancellations, it holds outside a measure-zero subset of structural equation models [13]. Hence, we adopt it here as a generic condition that enables consistent recovery of lagged causal structure in the presence of latent variables.
@@
3 CITS Outperforms Existing Models on Simulated Datasets
@@
Figure 2: Comparative study of inferring the Rolled graph. Inference of Rolled graph for five simulation settings (left to right): Linear Gaussian Models 1 and 2, Non-linear Non Gaussian Models 1 and 2, Continuous Time Recurrent Neural Network (CTRNN). Row 1: The ground truth for each simulation paradigm is graphically represented. The performances of the five methods Granger Causality 1 (GC1), Granger Causality 2 (GC2), Peter-Clark (PC), Time-Aware PC (TPC) and CITS, are shown in terms of three metrics (right column): 11-1 - False Positive Rate (IFPR) (green), True Positive Rate (TPR) (orange) and Combined Score (CS) (purple). Row 2 shows the performance the three metrics for $alpha=0.05$ and noise level 1.01.01.0. The CS of the methods over varying noise levels in simulation $\eta=0.1,0.5,1.0,\ldots,3.5$, with significance level $\alpha=0.01,0.05,0.1$ are also demonstrated in rows 3-5 respectively.
@@
We compare the performance of CITS, Pairwise Granger Causality (GC1), Multivariate Granger Causality (GC2), naive application of PC algorithm (PC), and Time-Aware PC algorithm (TPC), to recover the ground truth causal relations in simulated datasets (see Supplementary S1). We use simulated datasets from a variety of time series models ranging from linear to non-linear models, with and without common causes, and consider both the Gaussian and non-Gaussian noise settings (See Supplementary S5). In the simulations, for each model, we obtain 25 simulations of the entire time series each for different noise levels $\eta\in\{0.1,0.5,1,1.5,2,2.5,3,3.5\}$. All the time series simulated have $n=1000$ samples. We also used the level $\alpha$ of the conditional dependence test with $\alpha$ ranging in 0.01,0.050.01,0.050.01 , 0.05 and 0.10.10.1. The performance of the methods in recovering the ground truth causal relationships is summarized using the following three metrics: (1) Combined Score (CS), (2) True Positive Rate (TPR) and (3) 1 - False Positive Rate (IFPR) (see Methods at Section 5.2 for definitions). Conditional independence tests are based on partial correlations in Gaussian settings and the Hilbert-Schmidt criterion in non-Gaussian settings (see Section 5.2).
@@
In each of our simulation settings, there are 444 neurons and 161616 possible edges (including self-loops), leading to a total of 400400400 possible edges across 252525 simulations. CITS and TPC show superior performance in recovering the true Rolled graph at noise level $\eta=1$ and thresholding parameter $\alpha=0.05$ (Fig. 2). Their advantage is also evident in terms of Combined Score (CS) across varying noise levels $\eta$ and significance levels $\alpha$ in each simulation setting (Fig. 2). CITS consistently shows the strongest performance across all simulation settings, with TPC closely following. In the Linear Gaussian 1 scenario, where the true graph has three converging edges, CITS and TPC both achieve a TPR of 100% with high Combined Scores (97.2% and 96.6%, respectively), while GC1, despite having an IFPR of 100%, detects only one of the three true edges (TPR = 33.3%). In Linear Gaussian 2, which includes both a common cause and a common effect, CITS again leads with a CS of 99%, followed by TPC (41.7%) and GC2 (39%). Here, GC2’s performance is hindered by multicollinearity in regression, and PC shows poor sensitivity with a TPR of just 6%. In the Non-linear Non-Gaussian 1 scenario, CITS achieves a CS of 99.4%, outperforming all methods, while TPC follows closely (93.3%). GC2 detects all true edges (TPR = 100%) but introduces slightly more false positives. In the Non-linear Non-Gaussian 2 case, CITS again attains perfect scores across metrics (TPR = 100%, IFPR = 100%, CS = 100%). Finally, in the CTRNN setting with self-loops, CITS and TPC are the only methods that reliably detect recurrent structures, achieving TPRs of 88% and 85.7%, respectively. TPC achieves the highest CS (81.3%), just ahead of CITS (77.8%). Overall, CITS proves to be the most robust method across both linear and nonlinear settings, maintaining a high detection rate while minimizing false positives.
[…]
@@
The inferred edge weights and their directional nature (positive/increasing or negative/decreasing) at noise level $\eta=1$ and threshold $\alpha=0.05$ reveal that CITS most accurately recovers both the magnitude and sign of causal influences across scenarios (Fig. 3). Across all simulation settings, the edge weights inferred by CITS closely match the true values in both magnitude and sign. In the Linear Gaussian 1 scenario, the median estimates for 131$\rightarrow$ 31 → 3, 232$\rightarrow$ 32 → 3, and 343$\rightarrow$ 43 → 4 are nearly identical to the ground truth values (2, –1, and 2), with narrow ranges across trials and consistent sign recovery. Similarly, in Linear Gaussian 2, all four edges—including those involving common causes and effects—are accurately estimated, with medians near 2 and 1, and signs matching the true positive directions in every case.
In the Non-linear Non-Gaussian scenarios, despite the presence of sinusoidal dependencies and wider value ranges, CITS maintains high fidelity in edge weight estimation. Median values for connections such as 131$\rightarrow$ 31 → 3 and 343$\rightarrow$ 43 → 4 remain tightly centered around 3 to 3.5, while signs consistently reflect the correct increasing or decreasing influences. In the CTRNN setting, CITS detects self-loops with moderate strength (e.g., 0.73 to 0.94) and identifies strong feedforward edges such as 131$\rightarrow$ 31 → 3 and 232$\rightarrow$ 32 → 3 with high estimated weights (around 12). In all scenarios, the directional nature of each edge is recovered correctly across trials, highlighting the robustness of CITS to structural complexity and nonlinearity.
@@ 4 CITS Reveals Statistically Causal Neural Circuitry in the Mouse Brain During Visual Tasks
In neuroscience, the term Functional Connectivity (FC) refers to the network of interactions between individual units of the brain, such as neurons or brain regions, with respect to their activity over time [25, 2]. The main purpose of identifying the FC is to gain an understanding of how neurons work together to create brain function. The FC can be represented as a graph, where nodes denote neurons and edges denote a stochastic relationship between the activities of connected neurons. These edges can be undirected, indicating a stochastic association, whence the FC is termed as Associative FC (AFC). Alternatively, the edges can be directed and represent a statistically causal relationship between the activities of neurons, whence the FC is termed as Causal FC (CFC). Finding the CFC is expected to facilitate the inference of the governing neural interaction pathways essential for brain functioning and behavior [26, 2], and serves as a promising biomarker for neuro-psychiatric diseases [6]. The CFC is represented by a directed graph whose nodes are the neuron labels, and has an edge from neuron $u\rightarrow v$ if the activity of neuron $u$ at time $t$ has a statistically causal influence on the activity of neuron $v$ at a later time $t^{\prime}$ [1, 2, 12]. In terms of the framework described in Section 2.1, the CFC can be represented by the Rolled Graph of statistically causal neural interaction in the neural time series. Thereby, the CFC can be inferred by CITS algorithm (See Definition 1).In this section, we show how CITS reveals putative CFCs between neurons in the mouse brain during visual tasks. To do so, we analyzed high-resolution, large-scale electrophysiological recordings from the Allen Institute’s Visual Coding Neuropixels dataset in the Allen Brain Observatory [27]. The dataset includes sorted spike trains recorded simultaneously across six cortical visual areas, hippocampus, thalamus, and other adjacent structures of the mouse brain, while being presented with different types of visual stimuli. For our CFC analysis we chose responses to natural scenes, static gratings andGabor patches. We omitted responses to full-field flash stimuli from the analysis because such stimuli can evoke widespread, near-synchronous responses across the brain, leading to strong, stimulus-locked common inputs that inflate statistical dependencies and may masquerade as direct causal interactions. Prior work highlights how such extrinsic correlations, driven by the stimulus ensemble, can overwhelm intrinsic connectivity patterns [28, 29]. The visual stimuli are repeatedly presented to the mice and the data is recorded using the Neuropixels technology, consisting of multiple electrodes inserted into the brain allowing real-time recording from hundreds of neurons across brain different brain regions (For more details, see in Methods Section 5.1.)
4.1 Comparison of CITS and Other FC Methods.

### v2

# CITS: Nonparametric Statistical Causal Modeling for High‑Resolution Time Series

@@ Abstract
Identifying causal interactions in complex dynamical systems is a fundamental challenge across the computational sciences. Existing functional connectivity methods capture correlations but not causation. While addressing directionality, popular causal inference tools such as Granger causality and the Peter–Clark algorithm rely on restrictive assumptions that limit their applicability to high-resolution time-series data, such as the large-scale recordings now standard in neuroscience. Here, we introduce CITS (Causal Inference in Time Series), a nonparametric framework for inferring statistically causal structure from multivariate time series. CITS models dynamics using a structural causal model of arbitrary Markov order and statistical tests for lagged conditional independence. We prove consistency under mild assumptions and demonstrate superior accuracy over state-of-the-art baselines across simulated linear, nonlinear, and recurrent neural network benchmarks. Applying CITS to large-scale neuronal recordings from the mouse visual cortex, thalamus, and hippocampus, we uncover stimulus-specific causal pathways and inter-regional hierarchies that align with known anatomy while revealing new functional insights. We further highlight CITS ability in accurately identifying conditional dependencies within small inferred neuronal motifs. These results establish CITS as a theoretically grounded and empirically validated method for discovering interpretable statistically causal networks in neural time series. Beyond neuroscience, the framework is broadly applicable to causal discovery in complex temporal systems across domains.
@@ 1 Introduction
We evaluate CITS on a suite of synthetic time series benchmarks, including linear Gaussian, nonlinear non-Gaussian, and continuous-time recurrent neural networks. In each setting, CITS consistently recovers the correct causal structure more accurately than classical methods such as Granger Causality, the PC algorithm, and its time-aware extension (TPC), particularly in the presence of nonlinear dependencies and both source and sink in causal effects. We demonstrate the method’s practical utility in neuroscience to obtain functional neural circuitry under different visual stimulus conditions by applying CITS to high resolution electrophysiology recordings from mouse visual brain. We also uncovered the neural signal interactions within small neuronal motifs in the inferred pathways. The resulting causal graphs and their region-to-region anatomical distribution reveal biologically interpretable, stimulus-specific interactions among cortical, thalamic, and hippocampal regions, highlighting the potential of CITS to uncover directed neural interactions from large-scale electrophysiological time series.
2 CITS: A Statistically Causal Inference Framework for High-Resolution Time Series Data
@@
Figure 1: Inference of the Unrolled Directed Acyclic Graph (DAG) for Neural Time Series and its Rolled Graph. (A) Example Markovian Structural Causal Model interactions with $\tau=2$. (B) Formation of time-windowed samples for each $(v,t)$, where $0\leq t\leq 2\tau$. (C) The invariant unrolled DAG extends from time $t-\tau$, with edges projecting into it originating from at most time $t-2\tau$, motivating conditional dependence tests within a $2\tau$ window.
@@ 2.1 Markovian Structural Causal Model for Time Series
The graph $G$ associated with the SCM is constructed by creating one vertex for each $(v,t)$, where $v=1,\ldots,p$ and $t=1,\ldots,\tau+1$. A directed edge is drawn from each parent $(u,s)\in\text{pa}(v,\tau+1)$ to the node $(v,\tau+1)$, reflecting the functional dependence in (1). For notational simplicity, we may occasionally refer to the vertex $(v,t)$ as $X_{v,t}$. We assume that $G$ is acyclic, in which case it is a Directed Acyclic Graph (DAG). An illustrative example is shown in Fig. 1A.
@@
The graph $G$, constructed over variables indexed by $(v,t)$ with edges from $X_{u,t_{1}}$ to $X_{v,t_{2}}$ representing causal influence across time, is often referred to as the unrolled DAG. This representation explicitly captures the temporal evolution of causal dependencies over the sequence of observations. For interpretability and analysis, the unrolled DAG is often summarized into a rolled graph, also known as a summary causal graph [41]. The rolled graph $G_{R}$ is a directed graph over the $p$ components (variables), with an edge $u\to v$ if variable $X_{u,t}$ has a causal effect on $X_{v,t^{\prime}}$ at some later (or same) time $t^{\prime}\geq t$.
@@ 2.3.1 Oracle Version: With Conditional Independence Access
The properties described in Section 2.2 (see Lemma 1) motivate a two-level strategy for identifying $\text{pa}(v,t)$ using a conditional independence oracle. First, for each $u=1,\ldots,p$ and $s\in{t-\tau,\ldots,t}$, we test whether $X_{u,s}$ satisfies property (3). Second, for each such pair $(u,s)$, we search over subsets $S$ of $(d,r):d=1,\ldots,p;\ r=t-2\tau,\ldots,t$} to determine if conditional independence holds for at least one such $S$.
@@
6    if $X_{u,s}\perp\!\!\!\!\perp X_{v,2\tau+1}\penalty 10000\ |\penalty 10000\ \bm{X}_{S}$ then
@@ 2.4 Assigning Edge Weights in the Inferred Graphs
[…]
@@ Discussion.
This corollary follows from the temporal structure of the model. When all causal effects among observables are strictly time-lagged (i.e., from $X_{u,t-1}$ to $X_{v,t}$), with no concurrent (within-time-slice) effects, and the system is one-step Markovian, latent variables can induce at most contemporaneous dependencies within a single time step. Since CITS only tests for lagged conditional dependencies and conditions on past variables, it avoids inferring spurious lagged edges due to latent confounding under these assumptions. As long as the marginal distribution over observed variables is faithful to the induced causal graph, the consistency guarantees of Theorem 2.2 continue to apply. The assumption that the marginal distribution over observed variables is faithful to the marginal causal graph is standard in causal discovery. While faithfulness is not guaranteed under marginalization, since conditional independencies can arise through parameter cancellations, it holds outside a measure-zero subset of structural equation models [54]. Hence, we adopt it here as a generic condition that enables consistent recovery of lagged causal structure in the presence of latent variables.
@@
3 CITS Outperforms Existing Methods on Simulated Datasets
@@
Figure 2: Comparative study of inferring the Rolled graph. Inference of Rolled graph for five simulation settings (left to right): Linear Gaussian Models 1 and 2, Non-linear Non Gaussian Models 1 and 2, Continuous Time Recurrent Neural Network (CTRNN). Row 1: The ground truth for each simulation paradigm is graphically represented. The performances of the five methods Granger Causality 1 (GC1), Granger Causality 2 (GC2), Peter-Clark (PC), Time-Aware PC (TPC) and CITS, are shown in terms of three metrics (right column): 11- False Positive Rate (IFPR) (green), True Positive Rate (TPR) (orange) and Combined Score (CS) (purple). Row 2 shows the performance the three metrics for $alpha=0.05$ and noise level $1.0$. The CS of the methods over varying noise levels in simulation $\eta=0.1,0.5,1.0,\ldots,3.5$, with significance level $\alpha=0.01,0.05,0.1$ are also demonstrated in rows 3-5 respectively.
@@
We compare the performance of CITS, Pairwise Granger Causality (GC1), Multivariate Granger Causality (GC2), naive application of PC algorithm (PC), and Time-Aware PC algorithm (TPC), to recover the ground truth causal relations in simulated datasets (see Supplementary S1). We use simulated datasets from a variety of time series models ranging from linear to non-linear models, with and without common causes, and consider both the Gaussian and non-Gaussian noise settings (See Supplementary S5). In the simulations, for each model, we obtain 25 simulations of the entire time series each for different noise levels $\eta\in\{0.1,0.5,1,1.5,2,2.5,3,3.5\}$. All the time series simulated have $n=1000$ samples. We also used the level $\alpha$ of the conditional dependence test with $\alpha$ ranging in $0.01,0.05$ and $0.1$. The performance of the methods in recovering the ground truth causal relationships is summarized using the following three metrics: (1) Combined Score (CS), (2) True Positive Rate (TPR) and (3) 1 - False Positive Rate (IFPR) (see Methods at Section 5.2 for definitions). Conditional independence tests are based on partial correlations in Gaussian settings and the Hilbert-Schmidt criterion in non-Gaussian settings (see Section 5.2).
@@
In each of our simulation settings, there are $4$ neurons and $16$ possible edges (including self-loops), leading to a total of $400$ possible edges across $25$ simulations. CITS and TPC show superior performance in recovering the true Rolled graph at noise level $\eta=1$ and thresholding parameter $\alpha=0.05$ (Fig. 2). Their advantage is also evident in terms of Combined Score (CS) across varying noise levels $\eta$ and significance levels $\alpha$ in each simulation setting (Fig. 2). CITS consistently shows the strongest performance across all simulation settings, with TPC closely following. In the Linear Gaussian 1 scenario, where the true graph has three converging edges, CITS and TPC both achieve a TPR of 100% with high Combined Scores (97.2% and 96.6%, respectively), while GC1, despite having an IFPR of 100%, detects only one of the three true edges (TPR = 33.3%). In Linear Gaussian 2, which includes both a common cause and a common effect, CITS again leads with a CS of 99%, followed by TPC (41.7%) and GC2 (39%). Here, GC2’s performance is hindered by multicollinearity in regression, and PC shows poor sensitivity with a TPR of just 6%. In the Non-linear Non-Gaussian 1 scenario, CITS achieves a CS of 99.4%, outperforming all methods, while TPC follows closely (93.3%). GC2 detects all true edges (TPR = 100%) but introduces slightly more false positives. In the Non-linear Non-Gaussian 2 case, CITS again attains perfect scores across metrics (TPR = 100%, IFPR = 100%, CS = 100%). Finally, in the CTRNN setting with self-loops, CITS and TPC are the only methods that reliably detect recurrent structures, achieving TPRs of 88% and 85.7%, respectively. TPC achieves the highest CS (81.3%), just ahead of CITS (77.8%). Overall, CITS proves to be the most robust method across both linear and nonlinear settings, maintaining a high detection rate while minimizing false positives.
We then performed a systematic comparison of the Combined Score for CITS and other methods across noise levels $\eta$ ranging from $0.1$ to $3.5$ and significance levels $\alpha=0.01,0.05,0.1$ (Fig. 2). In the Linear Gaussian 1 scenario, CITS has a CS of nearly $100\%$ across noise levels greater than $1.0$ and all signifiance levels considered, closely matching the parametric GC2 model as well as non-parametric TPC, which are followed by GC1 in performance and lastly PC. In the Linear Gaussian 2 scenario, the distinction is remarkable, where CITS has a CS of $\approx 100\%$ across all levels of simulation noise and significance level $\alpha$, however the next best model for this setting are TPC and GC2 with a CS of $\approx 50\%$, followed by GC1 and lastly PC. In the Non-linear Non-Gaussian 1 and 2 scenarios, CITS has the highest CS compared to other methods across levels of noise and $\alpha$. In the CTRNN scenario, the best CS achieved is lower compared to other simulation paradigms. However, CITS and TPC have better performance compared to the other methods for noise level $\eta\geq 0.5$ and all $\alpha$.
@@
The inferred edge weights and their directional nature (positive/increasing or negative/decreasing) at noise level $\eta=1$ and threshold $\alpha=0.05$ reveal that CITS most accurately recovers both the magnitude and sign of causal influences across scenarios (Fig. 3). Across all simulation settings, the edge weights inferred by CITS closely match the true values in both magnitude and sign. In the Linear Gaussian 1 scenario, the median estimates for $1\rightarrow 3$, $2\rightarrow 3$, and $3\rightarrow 4$ are nearly identical to the ground truth values (2, –1, and 2), with narrow ranges across trials and consistent sign recovery. Similarly, in Linear Gaussian 2, all four edges—including those involving common causes and effects—are accurately estimated, with medians near 2 and 1, and signs matching the true positive directions in every case.
In the Non-linear Non-Gaussian scenarios, despite the presence of sinusoidal dependencies and wider value ranges, CITS maintains high fidelity in edge weight estimation. Median values for connections such as $1\rightarrow 3$ and $3\rightarrow 4$ remain tightly centered around 3 to 3.5, while signs consistently reflect the correct increasing or decreasing influences. In the CTRNN setting, CITS detects self-loops with moderate strength (e.g., 0.73 to 0.94) and identifies strong feedforward edges such as $1\rightarrow 3$ and $2\rightarrow 3$ with high estimated weights (around 12). In all scenarios, the directional nature of each edge is recovered correctly across trials, highlighting the robustness of CITS to structural complexity and nonlinearity.
@@ 4 CITS Reveals Statistically Causal Neural Circuitry in the Mouse Brain During Visual Tasks
In neuroscience, the term Functional Connectivity (FC) refers to the network of interactions between individual units of the brain, such as neurons or brain regions, with respect to their activity over time [45, 7]. The main purpose of identifying the FC is to gain an understanding of how neurons work together to create brain function. The FC can be represented as a graph, where nodes denote neurons and edges denote a stochastic relationship between the activities of connected neurons. These edges can be undirected, indicating a stochastic association, whence the FC is termed as Associative FC (AFC). Alternatively, the edges can be directed and represent a statistically causal relationship between the activities of neurons, whence the FC is termed as Causal FC (CFC). Finding the CFC is expected to facilitate the inference of the governing neural interaction pathways essential for brain functioning and behavior [14, 7], and serves as a promising biomarker for neuro-psychiatric diseases [36]. The CFC is represented by a directed graph whose nodes are the neuron labels, and has an edge from neuron $u\rightarrow v$ if the activity of neuron $u$ at time $t$ has a statistically causal influence on the activity of neuron $v$ at a later time $t^{\prime}$ [44, 7, 8]. In terms of the framework described in Section 2.1, the CFC can be represented by the Rolled Graph of statistically causal neural interaction in the neural time series. Thereby, the CFC can be inferred by CITS algorithm (See Definition 1).In this section, we show how CITS reveals putative CFCs between neurons in the mouse brain during visual tasks. To do so, we analyzed high‑resolution, large‑scale electrophysiological recordings from the Allen Institute’s Visual Coding Neuropixels dataset in the Allen Brain Observatory [24]. The dataset includes sorted spike trains recorded simultaneously across six cortical visual areas, hippocampus, thalamus, and other adjacent structures of the mouse brain, while being presented with different types of visual stimuli. For our CFC analysis we chose responses to natural scenes, static gratings andGabor patches. We omitted responses to full-field flash stimuli from the analysis because such stimuli can evoke widespread, near-synchronous responses across the brain, leading to strong, stimulus-locked common inputs that inflate statistical dependencies and may masquerade as direct causal interactions. Prior work highlights how such extrinsic correlations, driven by the stimulus ensemble, can overwhelm intrinsic connectivity patterns [39, 38]. The visual stimuli are repeatedly presented to the mice and the data is recorded using the Neuropixels technology, consisting of multiple electrodes inserted into the brain allowing real-time recording from hundreds of neurons across brain different brain regions (For more details, see in Methods Section 5.1.)
4.1 Comparison of CITS and Other FC Methods.
@@
Figure 5: Examples of Neural Signal Dependencies within Inferred CFC Motifs. A. Common-Source neuronal motif identified by CITS. In the inferred CFC, for the motif of neuron 105 $\leftarrow$ 125 $\rightarrow$ 247, pairwise plots exhibit correlations between all pairs including spurious correlation between 105 and 247 (red) even though not directly connected in the motif (top row). As predicited, conditioned on the parent, 105 and 125’s signals are uncorrelated. B. Multi-Parent neuronal motif identified by CITS. In a more complex motif of 106 $\leftarrow$ 105 $\leftrightarrow$ 118 $\rightarrow$ 109, pairwise plots exhibit correlations between all pairs (rows 1-2). As predicted, conditioned on both parents, neurons 106 and 109 become uncorrelated. This highlights CITS ability in correctly identifying these nodes as being not directly connected.
4.2 Neural Signal Dependencies within Inferred CFC Motifs
To interpret and validate the inferred CFC, we examined whether the neural signals obey the conditional independence patterns predicted by the graph structure. A defining feature of CFC, and what distinguishes it from standard functional connectivity, is its ability to isolate direct interactions from indirect network effects. Theoretically, this relies on the directed Markov property: two nodes may appear highly correlated due to a shared ancestor, but this dependence is rendered spurious (statistically independent) once the activity of the parent nodes is controlled.
We tested this phenomena by analyzing specific motifs within the inferred network. As shown in Fig. 5, marginally correlated neuron pairs often lack a direct connection in the CFC graph, implying their relationship is indirect. For example, in a "common source" motif (Fig. 5A), Neurons 105 and 247 exhibit a strong pairwise correlation (Pearson $r\approx 0.435$) despite not being connected. Consistent with the inferred CFC structure, conditioning on their shared parent (Neuron 125) explains away this relationship, reducing the residual correlation to near zero ($r\approx 0.06-0.08$).
This validation extends to more complex topologies involving multiple common drivers. In the motif shown in Fig. 5B, Neurons 106 and 109 share two distinct parents (Neurons 102 and 118) but have no direct link between them. While their raw activities are highly correlated (r=0.523), the CFC model correctly predicts that this is an indirect association. Indeed, when both parent neurons are conditioned on simultaneously, the correlation between the children vanishes ($r\approx 0.068$). These examples highlight the ability of CITS in identifying associations driven by upstream propagation and thereby providing a more rigorous map of neural information flow.